An astronaut is repairing a portion of the space station. In doing so, he threw a 5.25-kg monkey wrench away from him at 1.20 m/s to the right. If he has a mass of 72 kg, with what speed and in what direction will he begin to move



1
Expert's answer
2021-10-13T09:07:06-0400

Apply momentum conservation. The initial momentum is 0 (the astronaut and wrench are motionless), then the momentum is


"0=m_av_a+m_wv_w,\\\space\\
v_a=v_w\frac{m_w}{m_a}=1.2\frac{2.25}{72}=0.0375\text{ m/s}."